OpenStudy (anonymous):

Solve using the quadratic formula. (PLS show steps) 3x^2+5x=8

OpenStudy (jdoe0001):

\(\bf 3x^2+5x=8\implies 3x^2+5x-8=0 \\ \quad \\ \textit{quadratic formula}\\ y={\color{blue}{ 3}}x^2{\color{red}{ +5}}x{\color{green}{ +8}} \qquad \qquad x= \cfrac{ - {\color{red}{ b}} \pm \sqrt { {\color{red}{ b}}^2 -4{\color{blue}{ a}}{\color{green}{ c}}}}{2{\color{blue}{ a}}}\) plug and chug